I want to know about the differntial of following function , if it is possible
 
x(2x)(3x)(4x)............

Grade:11

1 Answers

Ananya
6 years ago
Let equation =yTake log both sidesLog y=log (2x+3x +4x)=logy=Log2x +log3x +log4xDifferentiate with respect to x1/y dy/dx=1/2x+1/2x+1/4xDy/dx=y(2x+3x +4x)Then put value of y from question and get answer.It`s simple.